# Decoupling Continual Semantic Segmentation

### v1

@@ B.1 Description of Classes
In our experiments, we employ GPT-4 [32] and DeepSeek-V3 [12] to generate these phrases. An automated script is then used to filter and clean the outputs, ensuring the generated descriptions are both concise and semantically rich.
@@ B.2 Training Details
Simultaneously, we introduced Asymmetric Loss [39] as an auxiliary loss function Eq. A.3, specifically designed to optimize the prediction $\mathbf{r}$ and image category classification results. This loss function is particularly suitable for handling class imbalance in multi-label classification problems, effectively adjusting the model’s sensitivity to different categories. This dual-loss mechanism ensures that the model maintains category detection accuracy while generating high-quality location-aware prompts.
@@ Appendix C Dataset
Datasets. Our method was evaluated on two semantic segmentation datasets PASCAL VOC2012 [14] and ADE20K [70]. PASCAL VOC 2012 serves as a relatively straightforward benchmark containing 20 object classes plus a background class, containing 10,582 samples for training and 1,449 for validation, while ADE20K presents a more challenging scenario which is a large-scale semantic segmentation dataset with 150 annotated classes and one background class, containing 20,210 and 2,000 samples for training and validation.

### v2

@@ B.1 Description of Classes
In our experiments, we employ GPT-4 [openai2024gpt4technicalreport] and DeepSeek-V3 [deepseekai2025deepseekv3technicalreport] to generate these phrases. An automated script is then used to filter and clean the outputs, ensuring the generated descriptions are both concise and semantically rich.
@@ B.2 Training Details
Simultaneously, we introduced Asymmetric Loss [ridnik2021asymmetric] as an auxiliary loss function Eq. A.3, specifically designed to optimize the prediction $\mathbf{r}$ and image category classification results. This loss function is particularly suitable for handling class imbalance in multi-label classification problems, effectively adjusting the model’s sensitivity to different categories. This dual-loss mechanism ensures that the model maintains category detection accuracy while generating high-quality location-aware prompts.
@@ Appendix C Dataset
Datasets. Our method was evaluated on two semantic segmentation datasets PASCAL VOC2012 [everinghamPascalVisualObject2010] and ADE20K [zhouSceneParsingADE20K2017]. PASCAL VOC 2012 serves as a relatively straightforward benchmark containing 20 object classes plus a background class, containing 10,582 samples for training and 1,449 for validation, while ADE20K presents a more challenging scenario which is a large-scale semantic segmentation dataset with 150 annotated classes and one background class, containing 20,210 and 2,000 samples for training and validation.